# On the degrees of freedom count on singular phase space submanifolds

### v3

@@ 1 Introduction
Initially, this paper of mine was initiated by the fact that it has been recently noticed [1] that a metric $f(R)$ gravity, with $f(R)=R^{2}$ and no linear term, has no dynamical degrees of freedom when linearised around Minkowski. It is of course immediately obvious for the usual graviton since $f^{\prime}(0)=0$, however might seem rather surprising for the scalar degree of freedom. Nevertheless this is indeed true, in the sense of maximally using the emergent gauge freedom of that limit. I would like to show that it can easily be seen directly from the equations of motion, with no need of playing games with Lagrangians. It is very important especially because those games are by far not always very innocent.
Moreover, the scalar part of this ”strong coupling” behaviour is actually more interpretational than dynamical. Unlike for the tensor graviton, it is not due to full disappearance of the mode from the quadratic action. It is not even a literal strong coupling of losing the kinetic energy only. It rather comes from losing a constraint. On the other hand, the different number of modes reported from a covariant procedure à la Stückelberg [1] has nothing to do with whatever remnant symmetries in the equations one might imagine. In the way the procedure had been applied, it does radically change the model at hand because of fixing a gauge directly inside the action.
@@ 2 Simple examples
  • If the variables are not dynamically predictable, and the freedom can be parametrised by arbitrary functions of all spacetime coordinates including time, I call it gauge freedom; and I also say that the number of those independent functions is the number of pure gauge modes. A non-trivial aspect here is that very often this is not a unique prescription, and we then usually go for minimising the number of modes in the dynamical regime below.
@@ 2.1 Scalar fields
Note though that sometimes these accidental constraints for momenta might be perfectly compatible with possibility of finding exact solutions. For example, the model of
@@
obviously produces a wave equation for $\phi^{3}$. This invertible change of variables, from $\phi$ to $\frac{1}{3}\phi^{3}$, removes the singularities from the Hamiltonian analysis. Those then present themselves through the fact that the solutions in terms of initial variables $\phi(x)=\sqrt[3]{\mathrm{a\ wave}}$ do have singular derivatives at the nodes of the wave.
More precisely, the model (3) above gives precisely the same results, irrespective of whether the initial variable $\phi$ or the new variable $\phi^{3}$ is used in the action, only due to its extreme simplicity. The equations $\phi^{2}\Box\phi^{3}=0$ and $\Box\phi^{3}=0$ do have the same set of smooth solutions because any derivative of an identical zero is also zero. At the same time, this change of variables, despite being invertible, modifies the variational procedure at the locus of $\phi=0$ which comes in terms of an extra factor of $\phi^{2}$ by which the two equations differ from each other. In more complicated models it might change the physics. For example, it has been used for producing new mimetic-like solutions [3, 4]. In my opinion [5], it is a very problematic procedure, in the sense of producing an ill-defined number of degrees of freedom. However, it is also an interesting aspect of singular behaviours.
@@ 2.2 An interlude: the Proca field
In any case, the transverse components must satisfy a simple wave equation. If $m=0$, this is the only dynamical information, while the temporal and longitudinal sector has one constraint of $\bigtriangleup A_{0}=\partial_{i}\dot{A}_{i}$ and one gauge freedom. If $m\neq 0$, divergence of the equation (8) imposes vanishing of the 4-vector divergence, and we easily come to the three degrees of freedom of the equations (7).
@@ 2.3 The Stückelberg trick
All in all, the physical vector $A_{\mu}+\partial_{\mu}\phi$ has got absolutely the same dynamics as the vector of the initial non-covariant model. The new scalar equation is simply about its divergencelessness. As the paper [1] correctly says referring to the classical review on massive gravity [6], the Stückelberg trick is not a reparametrisation but rather an introduction of a new variable. This is true. The catch is in the construction of a gauge freedom which leaves the invariant quantities the same as before the trick. Why it gave incorrect results in that paper [1] is because it had been used in an incorrect way.
Namely, instead of introducing a new variable, they did a reparametrisation of
@@
This reparametrisation is a bit redundant. One can shift the variables by $\partial_{\mu}\chi$ and $-\chi$ as long as the change satisfies an equation of $\Box\chi=0$. In case of separating a 3D-longitudinal mode, we fix such redundancy by spatial boundary conditions, but we treat the time differently. It can also be thought of as an incomplete gauge fixing of the model defined by the trick (10). This is all right at the level of equations of motion. However, a gauge fixing right inside the action generically changes the model.
@@
$\delta S=\int d^{4}x\left(\left(\partial_{\nu}F^{\nu\mu}+m^{2}{A^{T}}^{\mu}\right)\delta A^{T}_{\mu}-(\Box\phi)\delta\phi\right).$
Only the divergenceless part of the coefficient in front of $\delta A^{T}_{\mu}$ must vanish, but it automatically has this property, and this already corresponds to the three degrees of freedom of a massive vector field. The second term in the variation is one more, scalar degree of freedom. That’s how a gauge fixing inside the action or a reparametrisation with derivatives can drastically change a dynamical model.
One can go this way as far as (s)he wants. Even with a scalar theory, we can do a trick of $\phi\longrightarrow\phi+\Box\psi$. The new variables enjoy the gauge symmetry of $\phi\longrightarrow\phi+\Box\chi$ and $\psi\longrightarrow\psi-\chi$. If the initial model had an equation ${\mathcal{F}}[\phi]=0$, then the new one gets ${\mathcal{F}}[\phi+\Box\psi]=0$ and $\Box{\mathcal{F}}[\phi+\Box\psi]=0$. We’ve got a higher derivative model which is still equivalent to the initial one. Indeed, the first equation is precisely the same equation as it used to be, just acting on the gauge invariant variable, while the second equation follows from the first one, though in itself it has many more solutions. Had I taken a crazy idea of fixing the $\phi=0$ gauge right inside the action, it would have given me only the second equation, $\Box{\mathcal{F}}[\Box\psi]=0$, which is higher order even for the gauge invariant information, due to the variational prescription having been modified. Basically, the Lorenz-type gauges inside an action [1] commit the same crime. More elaborate examples of this very strange procedure can, for example, be found [^1] Just to be clear, I do not recommend taking it seriously. in Ref. [8].
@@ 2.4 Proca field with a variable mass
To a certain extent, the ugly amount of gauge freedom is an artefact of the linear approximation. Indeed, if we restrict ourselves to solutions with $\psi=0$ everywhere, then formally we simply have a massless vector field $\delta A_{\mu}$ with a gauge condition of $\delta A_{1}=0$ on top. This gauge fixing being incomplete leaves the remnant symmetry parametrised by $\chi(t,y,z)$. The problem is that it is not just a gauge choice but a requirement imposed by equations. Going to higher orders, we easily see that
@@
Any hope for improving the dynamical properties in such situations cannot then lie in taking higher orders of perturbation theory. The very model must then be changed. It’s a possible option in the spirit of effective field theory [10]. The model at hand might simply be some unsuccessful approximation to a more fundamental action. However, then the first question is whether a healthy and viable model exists at all, in a given class of theories. And even if the answer is affirmative, if the singular loci of the initial approximation were found in the physically interesting regions, then how can we ever trust anything we have seen there. Naively estimating the strong coupling scales [11] is just a possible way to close our eyes at the mathematical problem, for some time at least. Moreover, philosophically it is based on an idea of a kinetic term getting a vanishing coefficient, therefore producing an infinite factor in an interaction term upon canonical normalisation. It cannot always work like that in case of an accidental gauge symmetry though.
@@
The locus of $\psi=0$ corresponds to a massless vector field. Initial data of $\psi(0,\overrightarrow{x})=\dot{\psi}(0,\overrightarrow{x})=0$ are a separate world with gauge freedom in the full equations. It would be so even if we took a non-Abelian field which has its own equations non-linear. Going across this world is generically not smooth at all.
@@ 2.5 Non-linear self-interacting Proca
The main point for us now is that generically, away from $V^{\prime}=0$, we have no gauge freedom, one constrained mode $A_{0}$ and three dynamical modes $A_{i}$. Possible $V^{\prime}=0$ loci of non-vanishing $A_{\mu}$ do have a slightly more complicated structure of gauge freedom breaking, in the perturbations along the background field. Indeed, the quadratic Lagrangian around it gets a term of $2V^{\prime\prime}(A_{\mu}\delta A^{\mu})^{2}$, with the $4V^{\prime\prime}A^{\nu}A_{\mu}\cdot\delta A^{\mu}$ term in the equation of motion. For a purely spatial background field, we get a story of incomplete gauge freedom similar to what we have seen in the previous example (12).
@@
Generically, the initial velocities of the temporal and longitudinal modes get constrained thus removing one dynamical degree of freedom. However, if the initial values are chosen such that $3A^{2}_{0}=A^{2}_{j}$, the velocity $\dot{A}_{0}$ drops off. Then, despite the fact that the variable $A_{0}$ is fully detemined by the spatial components, those still have yet another constraint among themselves. The first equation (19) can be solved for $A_{0}$ as a function of $A_{i}$ and $\dot{A}_{i}$, then the second equation (20) gives a new constraint on the purely spatial Cauchy data, as long as $03A^{2}_{0}-A^{2}_{j}=$ 0 at the initial time. This can be easier seen in the Hamiltonian language to which I will come later.
@@ 3 $R^{2}$ gravity around Minkowski
was discussed concerning its number of degrees of freedom [1]. Obviously, it always enjoys $f^{\prime\prime}\neq 0$. However, precisely the Minkowski space, and every other vacuum GR solution, or even an arbitrary space of $R=0$ are on its problematic phase space surface, for $f^{\prime}(0)=0$. The usual gravitons obviously disappear from the linearised theory. It’s not a strong coupling in the literal meaning which would make them constrained, it is rather an accidental gauge symmetry totally removing the vector and tensor modes from the physical sector. At the same time, a probably less expected claim [1] is that the scalar mode is no longer alive either.
@@
which is basically the new scalar equation of motion, instead of the usual $R=0$ vacuum result, and then the equation (22) describes the two remaining polarisations of the graviton. Obviously, the scalar part of equations (23) survives taking the linear limit around the Minkowski metric. However, even a stronger requirement of $\partial_{\mu}\partial_{\nu}R=0$ does so (22).
Intuitively, the last remark already shows that the scalar mode must also go away because local perturbations of the scalar curvature have been killed. In the paper [1], more technical explanations for this fact are given. However, they intensively use derivative substitutions into the action which actually change the theory at hand. Therefore, I describe everything what happens directly at the level of equations of motion, and also show how the derivative changes of variables modify the theory leading the Authors of Ref. [1] to contradictory results. Basically, it all has already been illustrated on simple examples in the previous section.

### v5

@@ 1 Introduction
Initially, this paper of mine was initiated by the fact that it has been recently noticed [1] that a metric $f(R)$ gravity, with $f(R)=R^{2}$ and no linear term, has no dynamical degrees of freedom when linearised around Minkowski. It is of course immediately obvious for the usual graviton since $f^{\prime}(0)=0$, however might seem rather surprising for the scalar degree of freedom. Nonetheless, this is indeed true, in the sense of maximally using the emergent gauge freedom of that limit. I would like to show that it can easily be seen directly from the equations of motion, with no need of playing games with Lagrangians. It is very important, especially because those games are by far not always very innocent.
Moreover, the scalar part of this ”strong coupling” behaviour is actually more interpretational than dynamical. Unlike for the tensor graviton, it is not due to full disappearance of the mode from the quadratic action. It is not even a literal strong coupling of losing the kinetic energy only. It rather comes from losing a constraint. On the other hand, the same paper [1] also reported a different number of dynamical modes, namely the scalar degree of freedom having survived the linearisation procedure, via a covariant trick à la Stückelberg. This is another important topic to be discussed because the problem was due to an incorrect realisation of the trick. In the way it had been applied, it does radically change the model at hand because of fixing a gauge directly inside the action.
@@ 2 Simple examples
  • If the variables are not predictable at all, and the freedom can be parametrised by arbitrary functions of all spacetime coordinates including time, I call it gauge freedom; and I also say that the number of those independent functions is the number of pure gauge modes. A non-trivial aspect here is that very often this is not a unique prescription, which particular variable to call pure gauge, and we then usually go for minimising the number of modes in the dynamical regime below.
@@
Let me also stress it once more that I take the constrained modes as physical. It goes pretty much against the commonly used jargon of calling only dynamical ones physical. One might say that it is more of a philosophical question. However, taken at the face value, the common lore leads to proclaiming that Coulomb potentials in electrodynamics or Newtonian potentials in gravity are somehow unphysical. It does not seem reasonable. From the formal perspective of an abstract dynamical model, a pure gauge quantity is something that cannot be predicted at all from the equations of motion, and therefore unphysical for the given theory, while a constrained mode is fixed by the equations, with the only difference from dynamical modes in the need of Cauchy data or lack of that, and everything which follows from the equations I call physical.
One of the brightest examples of the opposite attitude can be found in a very recent paper [^1] Let me also use the occasion of mentioning this paper to admit that I totally do not consider the asymptotic symmetries. It is a very beautiful and interesting topic [3] on its own, however for now I prefer to think only in local terms. Note that, for formally deriving the equations of motion, it is enough to demand stationarity of the action with respect to perturbations of strictly finite support, and then no boundary terms are ever needed. [3] when they (correctly) say that ”it is not always straightforward to separate first-class constraints (that generate symmetries) from second-class constraints (that eliminate redundant fields)”. The catch is in the parentheses. Assuming that all first-class constraints generate (gauge) symmetries is about the extended Hamiltonian framework which I will discuss later. However, claiming that the second-class ones simply eliminate redundant fields means totally neglecting part of the information which is there in the equations of motion. If we accept that all what follows from equations is physical, then these fields are not redundant, they do carry some physics. At the same time, the pure gauge ones look much more redundant, in that they do not influence anything predictable.
@@ 2.1 Scalar fields
All in all, this is a very bad model (1) in that its Cauchy problem and therefore the number of degrees of freedom are ill-defined. It is not only about the approximation of quadratic action (i.e. linearised theory) around the problematic background of $\phi=0$. Generically, the set of possible solutions corresponds to two dynamical degrees of freedom, however it also has the singular set which contains $\phi\equiv 0$ and absolutely arbitrary $\chi$ as solutions of the full non-linear equations (2). We might treat it as a reasonable effective field theory only as long as we somehow can always stay far away from the zero value of the field $\phi$.
Note though that sometimes these accidental constraints for momenta might be perfectly compatible with regular structure of solutions. For example, a model of
@@
obviously produces a wave equation for $\phi^{3}$. If we attempted a Hamiltonian picture for the model (3), it would get an unwanted constraint for the momentum $p$ at the locus of $\phi=0$, though with a singular term $\frac{p^{2}}{2\phi^{4}}$ in the Hamiltonian which can easily be shown to go through a finite value at a node of the wave when $\phi(x)=\sqrt[3]{\mathrm{a\ wave}}$. The invertible change of variables, from $\phi$ to $\frac{1}{3}\phi^{3}$, removes the singularities from the Hamiltonian analysis. Those then present themselves through the fact that the solutions in terms of initial variables $\phi(x)=\sqrt[3]{\mathrm{a\ wave}}$ do have singular derivatives at the nodes of the wave.
2.2 Invertibility of a change of variables
It is important to stress that the model (3) above gives precisely the same results, irrespective of whether the initial variable $\phi$ or the new variable $\phi^{3}$ is used in the action, only due to its extreme simplicity. The equations
$\phi^{2}\Box\phi^{3}=0\quad\mathrm{and}\quad\Box\phi^{3}=0$
do practically have the same set of solutions because any derivative of an identical zero is also zero, so that an option of $\phi^{2}=0$ does not add anything new to the case of $\Box\phi^{3}=0$. At the same time, this change of variables $\phi\leftrightarrow\phi^{3}$, despite being invertible (in the realm of continuous functions), modifies the variational procedure at the locus of $\phi=0$ which comes in terms of an extra factor of $\phi^{2}$ by which the two equations differ from each other.
In more complicated models it might change the physics. For example, such changes of variables have been used for producing new mimetic-like solutions [4, 5]. Namely, they represent the physical metric as a disformal transformation of another one
$g_{\mu\nu}=C(\phi,{\tilde{X}})\cdot{\tilde{g}}_{\mu\nu}+D(\phi,{\tilde{X}})\cdot(\partial_{\mu}\phi)\partial_{\nu}\phi,\qquad{\tilde{X}}\equiv{\tilde{g}}^{\mu\nu}(\partial_{\mu}\phi)\partial_{\nu}\phi$
which is then treated as a new fundamental variable together with the auxiliary scalar field. One can easily see [6] that the metric transformation is invertible if and only if a simple scalar relation $X=X({\tilde{X}})$ is. Then it is not hard to show [7] that invertible cases do not produce anything new and can be called ”veiled general relativity (GR)”, while otherwise a particular mode of the physical metric gets represented by the scalar field only, and due to the extra derivatives in the action functional, it yields more general equations of mimetic gravity [8, 9].
The classical mimetic gravity [8] is about $X\equiv 1$, and therefore the transformation is never invertible. The new idea of the papers [4, 5] was in using a transformation with $\frac{dX}{d{\tilde{X}}}$ vanishing only for some particular value of $\tilde{X}$. At the price of an ill-defined number of degrees of freedom, one can get [4, 5] the mimetic type solutions on top of the pure GR for these special configurations of the scalar field. This is an example of one-to-one changes of variables being able to modify the model. With the derivative vanishing at an isolated point, the function $X(\tilde{X})$ can still be bijective and invertible in the class of continuous functions, but not continuously differentiable ones.
In my opinion [6], it is a very problematic procedure, in the sense of producing an ill-posed Cauchy problem. However, it is definitely an interesting and non-trivial aspect of singular behaviours. Note also that effects of various higher-derivative generalisations of such changes of variables inside the action are being actively studied nowadays [10].
@@ 2.3 An interlude: the Proca field
In any case, the transverse components must satisfy a simple wave equation. If $m=0$, this is the only dynamical information, while the temporal and longitudinal sector has one constraint of $\bigtriangleup A_{0}=\partial_{i}\dot{A}_{i}$ and one gauge freedom. If $m\neq 0$, divergence of the equation (8) imposes vanishing of the 4-vector divergence, and we easily come to the three dynamical degrees of freedom of the equations (7).
@@ 2.4 The Stückelberg trick
All in all, the physical vector $A_{\mu}+\partial_{\mu}\phi$ has got absolutely the same dynamics as the vector of the initial non-covariant model. The new scalar equation is simply about its divergencelessness. As the paper [1] correctly says referring to the classical review on massive gravity [11], the Stückelberg trick is not a reparametrisation but rather an introduction of a new variable. This is true. The key is in the construction of a gauge freedom which leaves the invariant quantities the same as before the trick.
However, using this trick, they had got incorrect results [1] on the number of degrees of freedom, i.e. more modes than in reality, for example the scalar dynamical mode in the purely $R^{2}$ gravity linearised around Minkowski. The reason is that it had been used in an incorrect way. Namely, instead of introducing a new variable, they did a reparametrisation of
@@
This reparametrisation is a bit redundant. One can shift the variables by $\partial_{\mu}\chi$ and $-\chi$ as long as the change satisfies an equation of $\Box\chi=0$. In case of separating a 3D-longitudinal mode, we may fix such redundancy by spatial boundary conditions, but we always treat the time differently. It can also be thought of as an (incomplete) gauge fixing [1] of the model defined by the trick (10). This is all right at the level of equations of motion. However, this gauge fixing right inside the action drastically changes the model.
@@
$\delta S=\int d^{4}x\left(\left(\partial_{\nu}F^{\nu\mu}+m^{2}{A^{T}}^{\mu}\right)\delta A^{T}_{\mu}-m^{2}(\Box\phi)\delta\phi\right).$
Only the divergenceless part of the coefficient in front of $\delta A^{T}_{\mu}$ must vanish, but it automatically has this property, and this already corresponds to the three degrees of freedom of a massive vector field. The second term in the variation is one more, scalar degree of freedom. That’s how a gauge fixing inside the action, or a reparametrisation with derivatives, can drastically change a dynamical model.
One can produce dynamics out of nothing. Consider a trivial action functional of the Lagrangian density ${\mathcal{L}}=A_{\mu}A^{\mu}$ with neither dynamical degrees of freedom nor gauge freedom. All four modes are constrained, $A_{\mu}=0$. If we do a Stückelberg trick $A_{\mu}\longrightarrow A_{\mu}+\partial_{\mu}\phi$, an obvious gauge freedom appears. However, either in the equations of motion or in the action, it leads to the very same equation for the physical variable $A_{\mu}+\partial_{\mu}\phi=0$ as before. At the same time, if one goes for a reparametrisation, or a gauge-fixed Stückelberg trick, $A_{\mu}=A^{T}_{\mu}+\partial_{\mu}\phi$ with $\partial_{\mu}{A^{T}}^{\mu}\equiv 0$, then it is all right at the level of equations, but inside the action it leads to more general equations, $A^{T}_{\mu}=0$ and $\square\phi=0$, which describe more freedom than before, even for the physical variable $A^{T}_{\mu}+\partial_{\mu}\phi$.
One can go this way as far as (s)he wants. Even with a scalar theory, we can do a trick of $\phi\longrightarrow\phi+\Box\psi$. The new variables enjoy the gauge symmetry of $\phi\longrightarrow\phi+\Box\chi$ and $\psi\longrightarrow\psi-\chi$. If the initial model had an equation ${\mathcal{F}}[\phi]=0$, then the new one gets ${\mathcal{F}}[\phi+\Box\psi]=0$ and $\Box{\mathcal{F}}[\phi+\Box\psi]=0$. We’ve got a higher derivative model which is still equivalent to the initial one. Indeed, the first equation is precisely the same equation as it used to be, just acting on the gauge invariant variable, while the second equation follows from the first one, though in itself it enjoys many more solutions. Had I taken a crazy idea of fixing the $\phi=0$ gauge right inside the action, it would have given me only the second equation, $\Box{\mathcal{F}}[\Box\psi]=0$, which is higher order even for the gauge invariant information, due to the variational prescription having been modified. Basically, the Lorenz-type gauges inside an action [1] commit the same crime. More elaborate examples of this very strange procedure can, for example, be found [^2] Just to be clear, I do not recommend taking it seriously. in Ref. [12].
@@ 2.5 Proca field with a variable mass
To a certain extent, the ugly amount of gauge freedom is an artifact of the linear approximation. Indeed, if we restrict ourselves to solutions with $\psi=0$ everywhere, then formally we simply have a massless vector field $\delta A_{\mu}$ with a gauge condition of $\delta A_{1}=0$ on top. This gauge fixing being incomplete leaves the remnant symmetry parametrised by $\chi(t,y,z)$. The problem is that it is not just a gauge choice but a requirement imposed by equations. Going to higher orders, we easily see that
@@
Any hope for improving the dynamical properties in such situations cannot then lie in taking higher orders of perturbation theory. The very model must then be changed. It’s a possible option in the spirit of effective field theory [14]. The model at hand might simply be some unsuccessful approximation to a more fundamental action. However, then the first question is whether a healthy and viable model exists at all, in a given class of theories. And even if the answer is affirmative, if the singular loci of the initial approximation were found in the physically interesting regions, then how can we ever trust anything we have seen there. Naively estimating the strong coupling scales [15] is just a possible way to close our eyes at the mathematical problem, for some time at least. Moreover, philosophically, reliance on adding new interaction terms when the initial coupling was strong is based on an idea that all the relevant variables stayed physical. It can be applied to cases of a kinetic term getting a vanishing coefficient, therefore producing an infinite factor in an interaction term upon canonical normalisation, but it cannot work like that for an accidental gauge symmetry though.
@@
The locus of $\psi=0$ corresponds to a massless vector field. Initial data of $\psi(0,\overrightarrow{x})=\dot{\psi}(0,\overrightarrow{x})=0$ are a separate world with gauge freedom in the full equations. It would be so even if we took a non-Abelian field which has its own equations non-linear. Going across this world is generically not smooth at all. And precisely like it was with the scalar field model (1), the model (14) is sick at the full non-perturbative level, not only in linear approximations. Generically, all the variables are physical, with 4 dynamical modes and one constrained quantity ($A_{0}$), however one can also have solutions with $\psi\equiv 0$ and one totally unpredictable mode in the vector field.
@@ 2.6 Non-linear self-interacting Proca
The main point for us now is that generically, away from $V^{\prime}=0$, we have no gauge freedom, one constrained mode $A_{0}$ and three dynamical modes $A_{i}$. Possible $V^{\prime}=0$ loci of non-vanishing $A_{\mu}$ do have a slightly more complicated structure of symmetry breaking, in the perturbations along the background field. Indeed, the quadratic Lagrangian around it gets a term of $2V^{\prime\prime}(A_{\mu}\delta A^{\mu})^{2}$, with the $4V^{\prime\prime}A^{\nu}A_{\mu}\cdot\delta A^{\mu}$ term in the equation of motion. For a purely spatial background field, we get a story of incomplete gauge freedom similar to what we have seen in the previous example (12).
@@
Generically, the initial velocities of the temporal and longitudinal modes get constrained thus removing one dynamical degree of freedom. However, if the initial values are chosen such that $3A^{2}_{0}=A^{2}_{j}$, the velocity $\dot{A}_{0}$ drops off. Then, despite the fact that the variable $A_{0}$ is fully detemined by the spatial components, those still have yet another constraint among themselves. The first equation (19) can be solved for $A_{0}$ as a function of $A_{i}$ and $\dot{A}_{i}$, then the second equation (20) gives a new constraint on the purely spatial Cauchy data, as long as $03A^{2}_{0}-A^{2}_{j}=$ 0 at the initial time. Again, it is a locus in the configuration space at which the Cauchy problem gets ill-defined. This can be easier seen in the Hamiltonian language to which I will come later.
@@ 3 $R^{2}$ gravity around Minkowski
was discussed concerning its number of degrees of freedom [1]. Obviously, it always enjoys $f^{\prime\prime}\neq 0$. However, precisely the Minkowski space, and every other vacuum GR solution, or even an arbitrary space of $R=0$ are on its problematic phase space surface, for $f^{\prime}(0)=0$. The usual gravitons immediately disappear from the linearised theory. In the present paper I do call it strong coupling, even though it is not the usual strong coupling in the literal meaning of the word which would make them constrained; it is rather an accidental gauge symmetry totally removing the vector and tensor modes from the physical sector. At the same time, a probably less expected claim [1] is that the scalar dynamical mode is no longer alive either.
@@
which is basically the new scalar equation of motion, instead of the usual $R=0$ vacuum result, and then the equation (22) describes the two remaining polarisations of the graviton. Obviously, the trace part of equations (23) survives taking the linear limit around the Minkowski metric. However, even a stronger requirement of $\partial_{\mu}\partial_{\nu}R=0$ does so (22).
Intuitively, the last remark already shows that the scalar mode must also go away because local perturbations of the scalar curvature have been killed. In the paper [1], more technical explanations for this fact are given. However, they present two methods of analysis, one of which indeed gives this answer while another one still shows a scalar mode in the spectrum. The reason is that they intensively use derivative substitutions into the action which generically change the theory at hand. Therefore, I describe everything what happens directly at the level of equations of motion, and also show how the derivative changes of variables modify the theory leading the Authors of Ref. [1] to the contradictory results. Basically, it all has already been illustrated on simple examples in the previous section.